OpenStudy (anonymous):

Solve 3x - 9x + 7 - 3 = -8. -6 -2 2 6

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Get variables onto one side, and constants onto another side: $$3x - 9x + 7 - 3 = -8$$ $$3x - 9x + 4 = -8 $$ $$3x - 9x = -8 - 4$$ $$-6x = -12$$ Divide both sides by -6: $$\frac{-6x}{-6}= \frac{-12}{-6}$$ $$x= 2$$

OpenStudy (anonymous):

You can plug x into the equation to verify if it is the correct answer.

OpenStudy (anonymous):

jezz ur smart I wish I was smart like you

OpenStudy (anonymous):

The internet is at your fingertips, use it wise. I recommended you look at khan academy, he covers most high school topics.